Determine the \\(curl F\\) at the point (2, 0, 3) given that \\(F=xz i+(2x^2-y)j-yz^2 k\\).
1
Expert's answer
2020-11-17T16:21:59-0500
"\text{curl}\vec{F}=\nabla\times\vec{F}=\begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 \dfrac{\partial}{\partial x} & \dfrac{\partial}{\partial y} & \dfrac{\partial}{\partial z} \\
xz & 2x^2-y & -yz^2
\end{vmatrix}="


"=\vec{i}(-z^2-0)-\vec{j}(0-x)+\vec{k}(4x-0)="

"=-z^2\vec{i}+x\vec{j}+4x\vec{k}"

"M(2,0,3)"


"\text{curl}\vec{F}_M=-9\vec{i}+2\vec{j}+8\vec{k}"